OpenStudy (anonymous):

CORRECT ME ^o^ :D Which of the following is the slope between the two points (-3, 1) and (3, -6)? six sevenths -six sevenths seven sixths -seven sixths <---- My answer !!

OpenStudy (anonymous):

-7/6 <--

OpenStudy (austinl):

\(\LARGE slope=m=\frac{\Delta rise}{\Delta run}=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

-6-1/3-3=-7/6 ?

OpenStudy (austinl):

\(\displaystyle \frac{-6-1}{3-(-3)}=\frac{-7}{6}\)

OpenStudy (anonymous):

Im right :D

OpenStudy (loser66):

yup

OpenStudy (loser66):

@austinL he needs confimation, :)

OpenStudy (austinl):

Well.... I am not entirely inclined to confirm as this person is being a little creepy in the chats...

OpenStudy (anonymous):

@Loser66 thank you :D ^.^